A magazine has 40 pages. If 16 of the 40 pages are advertising, what percentage of the magazine is advertising?
1. 2 over 5%
2. 16%
3. 24%
4. 40%

Answer by checkley77(12569)

16/40=2/5=.4=40% ans.(4)